# CANDLE: Iterative Conceptualization and Instantiation Distillation from Large Language Models for Commonsense Reasoning

### v1

@@ Abstract
The sequential process of conceptualization and instantiation is essential to generalizable commonsense reasoning as it allows the application of existing knowledge to unfamiliar scenarios. However, existing works tend to undervalue the step of instantiation and heavily rely on pre-built concept taxonomies and human annotations to collect both types of knowledge, resulting in a lack of instantiated knowledge to complete reasoning, high cost, and limited scalability. To tackle these challenges, we introduce CANDLE, a distillation framework that iteratively performs contextualized conceptualization and instantiation over commonsense knowledge bases by instructing large language models to generate both types of knowledge with critic filtering. By applying CANDLE to ATOMIC, we construct a comprehensive knowledge base comprising six million conceptualizations and instantiated commonsense knowledge triples. Both types of knowledge are firmly rooted in the original ATOMIC dataset, and intrinsic evaluations demonstrate their exceptional quality and diversity. Empirical results indicate that distilling CANDLE on student models provides benefits across four downstream tasks. Our code, data, and models are publicly available at https://github.com/HKUST-KnowComp/CANDLE.
@@ 1 Introduction
Commonsense reasoning refers to the cognitive ability to make logical inferences and draw conclusions based on general knowledge and understanding of the world that is typically shared among individuals Davis (2014); Mueller (2014). However, a longstanding challenge is its generalizability, as it often necessitates applying knowledge and reasoning to novel situations and contexts beyond simple pattern recognition or memorizing all special cases Mortimer (1995); Banaji and Crowder (1989). One promising approach to address this is the chain of conceptualization Murphy (2004) and instantiation Anderson et al. (1976), which, akin to the process of conceptual induction and deduction in human reasoning Tenenbaum et al. (2011), involves conceptualizing instances derived from known commonsense knowledge and subsequently instantiating these concepts in new situations to obtain the knowledge required for downstream reasoning. For example, in Figure 1, one can first conceptualize exercising in the gym as a healthy lifestyle, and further instantiate it to go on a balanced diet. This process allows the derivation of a novel event, PersonX goes on a balanced diet, which may entail new commonsense knowledge when placed back in the original context. By possessing substantial knowledge to initiate the process of conceptualization and instantiation, one can extrapolate limited commonsense knowledge to a wide array of diverse scenarios.
@@
Yet, replicating this fundamental ability on machines remains challenging due to the absence of both types of knowledge in widely used CommonSense Knowledge Bases (CSKBs; Sap et al., 2019a; Speer et al., 2017; Fang et al., 2021b, a). Language models that are fine-tuned on these CSKBs are, therefore, unable to effectively utilize such a chain for reasoning during inference. To compensate, various methods have been proposed for entity-level Durme et al. (2009); Song et al. (2011, 2015); Gong et al. (2016); He et al. (2020); Peng et al. (2022a) and event-level Chen et al. (2020); He et al. (2022a); Wang et al. (2023b) conceptualizations by matching against concept taxonomies like Probase Wu et al. (2012) and WordNet Miller (1995). Meanwhile, Allaway et al. (2023) address instantiation through controllable generation. However, several limitations still persist.
@@
Firstly, the two-step chain of conceptualization and instantiation can significantly augment a CSKB by introducing a diverse range of instances not originally included, thereby expanding the CSKB’s semantic coverage. However, most existing works underestimate the importance of the second step, instantiation, by focusing solely on conceptualization and using the resulting abstract knowledge directly. Other studies that concentrate on instantiations either overlook the conceptualization step entirely or only retrieve instances from the original CSKB, failing to introduce novel entities and events. Secondly, most conceptualization methods heavily depend on matching instances with concepts in concept taxonomies, such as Probase and WordNet, which have a limited scope and lack contextual information. Consequently, the derived conceptualizations are constrained in scale by these taxonomies and are formulated without considering proper contextualization, necessitating further verification in the original context. Lastly, the chain of conceptualization and instantiation can easily bring more than two orders of magnitude of data on top of the original CSKB. However, current acquisition and verification methods for both steps heavily rely on human annotation, which can be extremely costly as the scale of the CSKB increases.
@@
By applying CANDLE to ATOMIC Sap et al. (2019a), we construct a large-scale knowledge base comprising 6.18 million conceptualizations and instantiations from two powerful LLMs, ChatGPT OpenAI (2022) and LLAMA2 Touvron et al. (2023). We demonstrate the intrinsic efficacy of CANDLE through automatic and human evaluations, showcasing the superior quality and diversity of knowledge generated by CANDLE (Section 5.1). We further show the extrinsic benefits of CANDLE by leveraging the generated knowledge as complementary training data to distill student models that yield improvements across four downstream tasks, including CSKB conceptualization, generative commonsense inference, and zero-shot commonsense question answering (Section 5.2).
@@ 2.1 Conceptualization and Instantiation
Conceptualization aims to abstract a set of entities or events into a general concept, thereby forming abstract commonsense knowledge within its original context Murphy (2004). Subsequently, instantiation ground the derived concept into other instances and events to introduce new commonsense knowledge. Existing works primarily focused on entity-level conceptualization Durme et al. (2009); Song et al. (2011, 2015); Liu et al. (2022); Peng et al. (2022a), with He et al. (2022a) pioneering the construction of an event conceptualization benchmark. They combined syntactic parsing, semantically heuristic matching, and human annotation to extract concepts for social events from WordNet Miller (1995) synsets and Probase Wu et al. (2012). Building upon this work, Wang et al. (2023b) proposed a semi-supervised framework for conceptualizing CSKBs and demonstrated that abstract knowledge can enhance commonsense inference modeling and commonsense question answering Wang et al. (2023a). Wang et al. (2023d) also extended the study by constructing an abstraction benchmark based on eventualities from ASER Hongming Zhang et al. (2022). Regarding instantiation, Allaway et al. (2023) introduced a controllable generative framework to identify valid instantiations for abstract knowledge automatically. However, none of the existing studies have fully completed the chain of conceptualization and instantiation, with each focusing on only one aspect. Human annotation is also frequently applied for data collection and verification, which is both expensive and limited in scalability. Additionally, the downstream benefits of instantiated commonsense knowledge have not been thoroughly explored, leaving a significant gap in improving commonsense reasoning models.
@@ 2.2 Commonsense Knowledge Distillation
Recent breakthroughs in LLMs OpenAI (2022, 2023) have led to numerous efforts in distilling commonsense knowledge into datasets for training performant student models. West et al. (2022); Sclar et al. (2022); Bhagavatula et al. (2023); West et al. (2023) followed the pipeline of symbolic knowledge distillation, which uses human-crafted prompts to extract specific types of knowledge from LLMs for training downstream models. He et al. (2022b) proposed an alternative that involves transferring distilled knowledge from a ranker to a retriever, resulting in a more consistent and robust commonsense generator. Chae et al. (2023) and Kim et al. (2023) focused on distilling conversational responses from LLMs to enhance dialogue agents with commonsense knowledge and high-quality rationales. In this paper, we share similar aspirations and propose a chain of distillation framework that sequentially obtains abstract and instantiated knowledge from powerful LLMs. Empirical results show that our framework offers more substantial downstream benefits than previous symbolic knowledge distillation methods.
@@ 3 Definitions and Datasets
We follow the definitions proposed by He et al. (2022a) and Wang et al. (2023b) to formulate conceptualization and instantiation. Denote the triples in the original CSKB as , where , , and are the set of heads, relations, and tails in the original CSKB. The objective of conceptualization is to form a conceptualized head event, denoted as , from the original head . This is achieved by linking a component to a concept , forming by replacing with . Consequently, abstract knowledge is formed by combining the conceptualized head event with the original relation and tail, represented by . In the next step, the goal of instantiation is to associate the concept with a new instance . This process enables the formation of new commonsense knowledge in the format of , where is obtained by replacing with . In this paper, we use ATOMIC Sap et al. (2019a) as the original CSKB , which contains 310K triples after dropping those with wildcards and 18,839 unique head events. AbstractATOMIC He et al. (2022a) is used as the source of instances for every head event .
@@ 4.1 Contextualized Conceptualization
Previous methods for collecting conceptualizations rely on heuristically matching instances against concepts from WordNet and Probase. However, they suffer from limited concept coverage, resulting in a lack of knowledge diversity after instantiation, and require additional verification to ensure that concept fits into the original context . To address both issues, we propose to utilize ChatGPT as a loose teacher to collect conceptualizations in a one-step inference manner. Following Brown et al. (2020) and West et al. (2022), we use a few-shot prompt to instruct ChatGPT:
@@
| <EX-INP><EX-OUT> |
| <EX-INP><EX-OUT> |
| <EX-INP> |
where <TASK-PROMPT> is a task instruction that explains how to conceptualize an event and <EX-INP><EX-OUT> are human authored examples of conceptualizations for events sampled from ATOMIC. For each example, are included in the input, and is the output. Finally, we provide the N input as <EX-INP> and ask ChatGPT to generate the corresponding conceptualization as <EX-OUT>. This ensures that ChatGPT not only learns the relationship between instances and their conceptualizations but also performs such abstraction in a contextualized manner, ensuring the plausibility of the generated conceptualization within the original context . Additionally, we include several hints in the prompt, such as numbering the examples, generating concepts instead of hypernyms, and keeping the generated responses concise. In this paper, we set N = 11 and obtain 20 conceptualizations for every event .
@@ 4.2 Contextualized Instantiation
After conceptualizing all events, we proceed to instantiate them by instructing an LLM once again. We use LLAMA2-13B as the loose teacher model for instantiation to reduce the significant cost associated with instantiating numerous conceptualizations. Our pilot study demonstrates that it can also generate high-quality instantiation knowledge. We employ a similar prompt as described in Section 4.1, with the modification of replacing <TASK-PROMPT> with the explanation of instantiating a conceptualized event and changing <EX-INP><EX-OUT> to human-authored examples of instantiations for abstract commonsense knowledge triples. are included in the input and is the expected output. By learning from these examples, LLAMA2 is expected to generate the corresponding instantiation (<EX-OUT>) based on the given abstract knowledge triple (<EX-INP>). We also set N = 11 and produce only one instantiation for each conceptualized event due to the significant amount of conceptualizations obtained in the previous step. Appendix A provides more details regarding the distillation process.
@@ 4.3 Iterating with Critic Filtering
Following West et al. (2022), we use critic filtering models to eliminate low-quality generations from LLMs. Specifically, we utilize a DeBERTa-v3-large conceptualization discriminator, provided by Wang et al. (2023b), and VERA-T5-xxl, provided by Liu et al. (2023), to evaluate the quality of the generated conceptualizations and instantiations, respectively. We set an empirical threshold value to serve as the cutoff point for discarding generations with scores below . In Section 5.1, we present evaluations conducted to determine the optimal value for . For all downstream applications, we set . Post-filtering, the instantiated triples can be reintroduced as the input for conceptualizations again as they continue to represent concrete commonsense knowledge. This iterative process of conceptualization and instantiation forms a loop, which enables continuously augmenting a CSKB. In this paper, we execute the loop only once, but multiple iterations hold the promise of significantly enhancing the CSKB’s knowledge coverage.
@@ 5.1 Distillation Evaluations
We present the statistics of CANDLE distillation on ATOMIC in Table 1, demonstrating that CANDLE surpasses other existing benchmarks in terms of scale and concept coverage. Even under a rigorous critic filtering threshold (), CANDLE maintains its leading position with the highest count of total and unique knowledge for both types. To assess the quality of the distilled knowledge, we recruit four expert annotators, who are graduate students with expertise in commonsense, to conduct human evaluations on the plausibility of the generated conceptualizations and instantiations. Specifically, we randomly select 3,000 abstract commonsense triples and 3,000 instantiated triples from the distilled knowledge set. The annotators are then asked to determine (1) the correctness of the conceptualizations and instantiations and (2) the plausibility of their formed triples. Only triples whose both parts are deemed plausible by all annotators are considered accepted. We analyze the ratio of accepted triples among the samples with different levels of critic filtering and report the results in Table 1. Our findings demonstrate that LLMs exhibit impressive abilities in conceptualization and instantiation, with initial plausibility rates of 82.6% and 77.9% for both types of knowledge, respectively. With critic filtering, we observe an expected improvement in plausibility by maximally 14.6% and 16.6%, underscoring the effectiveness of our filtering measures in ensuring the high quality of the distilled knowledge. Details about annotations are in Appendix D.
@@
The process of abstracting an event into highly diverse conceptualizations plays a crucial role in CANDLE. It is of significant importance because the greater the diversity of conceptualizations, the broader the knowledge coverage becomes upon instantiation. This, in turn, enhances the overall knowledge coverage within the distillation process. To examine the diversity of the top 10,000 popular distilled conceptualizations, we obtain their hypernyms by matching against Probase Wu et al. (2012) and present a visualization in Figure 3. It reveals that our distilled conceptualizations demonstrate a high level of diversity across various categories, forming a comprehensive and intricate knowledge base.
@@
| CANDLE Distilled (Ours) | RoBERTa-large 340M | 80.69 | 80.99 | 83.11 | 84.50 |
| DeBERTa-v3-large 435M | 80.97 | 81.14 | 83.64 | 84.64 | |
| LLAMA2 (LoRA Fine-tuned) 7B | 77.48 | 78.27 | 81.68 | 83.40 | |
| Mistral-v0.1 (LoRA Fine-tuned) 7B | 77.77 | 78.29 | 81.95 | 82.54 | |
| VERA-T5 (Fine-tuned) 5B | 77.54 | 78.03 | 82.79 | 83.61 | |

### v2

@@ Abstract
The sequential process of conceptualization and instantiation is essential to generalizable commonsense reasoning as it allows the application of existing knowledge to unfamiliar scenarios. However, existing works tend to undervalue the step of instantiation and heavily rely on pre-built concept taxonomies and human annotations to collect both types of knowledge, resulting in a lack of instantiated knowledge to complete reasoning, high cost, and limited scalability. To tackle these challenges, we introduce CANDLE (ConceptuAlization and INstantiation Distillation from Large Language ModEls), a distillation framework that iteratively performs contextualized conceptualization and instantiation over commonsense knowledge bases by instructing large language models to generate both types of knowledge with critic filtering. By applying CANDLE to ATOMIC Sap et al. (2019a), we construct a comprehensive knowledge base comprising six million conceptualizations and instantiated commonsense knowledge triples. Both types of knowledge are firmly rooted in the original ATOMIC dataset, and intrinsic evaluations demonstrate their exceptional quality and diversity. Empirical results indicate that distilling CANDLE on student models provides benefits across three downstream tasks111Our data and models are publicly available at https://github.com/HKUST-KnowComp/CANDLE..
@@ 1 Introduction
Commonsense reasoning refers to the cognitive ability to make logical inferences and draw conclusions based on general knowledge and understanding of the world that is typically shared among individuals Davis (2014); Mueller (2014). However, a longstanding challenge is generalizability, as commonsense reasoning often necessitates applying knowledge to novel situations beyond simple pattern recognition or memorizing all special cases Mortimer (1995); Banaji and Crowder (1989). One promising approach to address this is the chain of conceptualization Murphy (2004) and instantiation Anderson et al. (1976), which, akin to the process of conceptual induction and deduction in human reasoning Tenenbaum et al. (2011), involves conceptualizing instances derived from known commonsense knowledge and subsequently instantiating these concepts in new situations to obtain the knowledge required for downstream reasoning. For example, in Figure 1, one can first conceptualize enjoys exercising in the gym as a healthy lifestyle, and then further instantiate it to go on a balanced diet. This process allows for the derivation of a novel event, PersonX goes on a balanced diet, which may entail new commonsense knowledge when connected with the original event’s commonsense inferential tail. By possessing substantial knowledge to initiate the process of conceptualization and instantiation, one can extrapolate limited commonsense knowledge to a wide array of diverse scenarios.
@@
Yet, replicating this fundamental ability on machines remains challenging due to the absence of both types of knowledge in widely used CommonSense Knowledge Bases (CSKBs; Sap et al., 2019a; Speer et al., 2017; Fang et al., 2021b, a). Various methods compensating the lack of conceptualization ability of language models have been proposed for entity-level Durme et al. (2009); Song et al. (2011, 2015); Gong et al. (2016); He et al. (2020); Peng et al. (2022a) and event-level Chen et al. (2020); He et al. (2024); Wang et al. (2023b) conceptualizations by matching against concept taxonomies like Probase Wu et al. (2012) and WordNet Miller (1995). However, several limitations still persist.
@@
Firstly, despite the importance of both conceptualization and instantiation, most existing works underestimate the importance of the second step while focusing solely on conceptualization and using the resulting abstract knowledge directly. Other studies that concentrate on instantiations either overlook the conceptualization step entirely or only retrieve instances from the original CSKB, failing to introduce novel entities and events. Secondly, most conceptualization methods heavily depend on matching instances with concepts in concept taxonomies, such as Probase and WordNet, which have a limited scope and lack contextual information. Consequently, the derived conceptualizations are constrained in scale by these taxonomies and are formulated without considering proper contextualization, necessitating further verification in the original context. Lastly, the chain of conceptualization and instantiation can easily bring more than two orders of magnitude of data on top of the original CSKB. However, current acquisition and verification methods for both steps heavily rely on human annotation, which can be extremely costly as the scale of the CSKB increases.
@@
By applying CANDLE to ATOMIC Sap et al. (2019a), we construct a large-scale knowledge base comprising 6.18 million conceptualizations and instantiations from two powerful LLMs, ChatGPT OpenAI (2022) and LLAMA2 Touvron et al. (2023). We demonstrate the intrinsic efficacy of CANDLE through automatic and human evaluations, highlighting the ability to generate high-quality and diverse knowledge (Section 5.1). We further show the extrinsic benefits of CANDLE by leveraging the generated knowledge as complementary training data to distill student models that yield improvements across three downstream tasks, including CSKB conceptualization, generative commonsense inference, and zero-shot commonsense question answering (Section 5.2).
@@ 2.1 Conceptualization and Instantiation
Conceptualization aims to abstract a set of entities or events into a general concept, thereby forming abstract commonsense knowledge within its original context Murphy (2004). Subsequently, instantiation grounds the derived concept into other instances and events to introduce new commonsense knowledge. Existing works primarily focused on entity-level conceptualization Durme et al. (2009); Song et al. (2011, 2015); Liu et al. (2022); Peng et al. (2022a), with He et al. (2024) pioneering the construction of an event conceptualization benchmark by extracting concepts for social events from WordNet Miller (1995) synsets and Probase Wu et al. (2012). Wang et al. (2023b, a) further proposed a semi-supervised framework for conceptualizing CSKBs and demonstrated that abstract knowledge can enhance commonsense inference modeling and question answering. Wang et al. (2023d) constructed an abstraction benchmark based on eventualities from ASER Zhang et al. (2022). Regarding instantiation, Allaway et al. (2023) introduced a controllable generative framework to identify valid instantiations for abstract knowledge automatically. However, none of the existing studies have fully completed the chain of conceptualization and instantiation, with each focusing on only one aspect. Human annotation is also frequently applied for data collection and verification, which is both expensive and limited in scalability. Additionally, the downstream benefits of instantiated commonsense knowledge have not been thoroughly explored, leaving a significant gap in improving commonsense reasoning models.
@@ 2.2 Commonsense Knowledge Distillation
Recent breakthroughs in LLMs OpenAI (2022, 2023) have led to numerous efforts in distilling commonsense knowledge into datasets for training performant student models. West et al. (2022); Sclar et al. (2022); Bhagavatula et al. (2023); West et al. (2023) followed the pipeline of symbolic knowledge distillation, which uses human-crafted prompts to extract specific types of knowledge from LLMs for training downstream models. He et al. (2022) proposed to transfer distilled knowledge from a ranker to a retriever, resulting in a more robust commonsense generator. Chae et al. (2023) and Kim et al. (2023) focused on distilling conversational responses from LLMs to enhance dialogue agents with commonsense knowledge and high-quality rationales. In this paper, we share similar aspirations and propose a chain of distillation framework that sequentially obtains abstract and instantiated knowledge from powerful LLMs. Empirical results show that our framework offers more substantial downstream benefits than traditional symbolic knowledge distillation methods.
@@ 3 Definitions and Datasets
We follow the definitions proposed by He et al. (2024) and Wang et al. (2023b) to formulate conceptualization and instantiation. Denote the triples in the original CSKB as , where , , and are the set of heads, relations, and tails in the original CSKB. The objective of conceptualization is to form a conceptualized head event, denoted as , from the original head . This is achieved by linking a component to a concept , forming by replacing with . Consequently, abstract knowledge is formed by combining the conceptualized head event with the original relation and tail, represented by . In the next step, the goal of instantiation is to associate the concept with a new instance . This process enables the formation of new commonsense knowledge in the format of , where is obtained by replacing with . In this paper, we use ATOMIC Sap et al. (2019a) as the original CSKB , which contains 310K triples after dropping those with wildcards and 18,839 unique head events. AbstractATOMIC He et al. (2024) is used as the source of instances for every head event .
@@ 4.1 Contextualized Conceptualization
Previous methods for collecting conceptualizations rely on heuristically matching instances against concepts from WordNet and Probase. However, they suffer from limited concept coverage, resulting in a lack of knowledge diversity after instantiation, and require additional verification to ensure that concept fits into the original context . To address both issues, we propose to utilize ChatGPT as a loose teacher to collect conceptualizations in a one-step inference manner. To verify the feasibility of such choice instead of other open-source LLMs, we carry out a pilot study, in which we randomly select 1000 events and asked ChatGPT and LLAMA2-7B to generate their conceptualizations. The results of our expert evaluation show that ChatGPT has 98% plausible generations, while LLAMA2 only achieves 81%. Therefore, we choose ChatGPT as our core conceptualizer due to its exceptional performances. Following Brown et al. (2020) and West et al. (2022), we use a few-shot prompt to instruct ChatGPT:
@@
| <EX1-INP><EX1-OUT> |
| <EXN-1-INP><EXN-1-OUT> |
| <EXN-INP> |
where <TASK-PROMPT> is a task instruction that explains how to conceptualize an event and <EX1-INP><EX1-OUT> are human authored examples of conceptualizations for events sampled from ATOMIC. For each example, are included in the input, and is the output. Finally, we provide the N input as <EXN-INP> and ask ChatGPT to generate the corresponding conceptualization as <EXN-OUT>. This ensures that ChatGPT not only learns the relationship between instances and their conceptualizations but also performs such abstraction in a contextualized manner, ensuring the plausibility of the generated conceptualization within the original context . In this paper, we set = 6 and obtain = 20 conceptualizations for every event .
@@ 4.2 Contextualized Instantiation
After conceptualizing all events, we proceed to instantiate them by instructing an open-source LLM to reduce the cost as the scale of instantiation is = 20 times larger than that of conceptualization. Similarly, we carry out a round of pilot study to demonstrate the feasibility of employing LLAMA2. We ask both ChatGPT and LLAMA2 to instantiate 1,000 ChatGPT-generated conceptualizations, and find that both models are able to produce approximately 95% plausible instantiations with critic filtering. Considering the significant cost of using ChatGPT to generate 6.18 million conceptualizations, we decide to use LLAMA2-13B as our core instantiater (Case studies are shown in Appendix G). We employ a similar prompt as described in Section 4.1, with the modification of replacing <TASK-PROMPT> with the explanation of instantiating a conceptualized event and changing <EX1-INP><EX1-OUT> to human-authored examples of instantiations for abstract commonsense knowledge triples. are included in the input and is the expected output. By learning from these examples, LLAMA2 is expected to generate the corresponding instantiation (<EXN-OUT>) based on the given abstract knowledge triple (<EXN-INP>). We set = 11 and produce only one instantiation for each conceptualized event due to the significant amount of conceptualizations obtained in the previous step. Appendix A provides more details regarding the distillation process.
@@ 4.3 Iterating with Critic Filtering
Following West et al. (2022), we use critic filtering models to eliminate low-quality generations from LLMs. Specifically, we utilize a DeBERTa-v3-large conceptualization discriminator, provided by Wang et al. (2023b), and VERA-T5-xxl, provided by Liu et al. (2023), to evaluate the quality of the generated conceptualizations and instantiations, respectively. We set an empirical threshold value to serve as the cutoff point for discarding generations with scores below . In Section 5.1, we present evaluations conducted to determine the optimal value for . For all downstream applications, we set = 0.9. Post-filtering, the instantiated triples can be reintroduced as the input for conceptualizations again as they continue to represent concrete commonsense knowledge. This iterative process of conceptualization and instantiation forms a loop, which enables continuously augmenting a CSKB. In this paper, we execute the loop only once, but multiple iterations hold the promise of significantly enhancing the CSKB’s knowledge coverage.
5 Evaluations and Analysis
@@ 5.1 Distillation Evaluations
We present CANDLE distillation statistics based on ATOMIC in Table 1, showing its superiority in scale and concept coverage compared to other benchmarks. Even with a strict critic filtering threshold ( = 0.9), CANDLE maintains its leading position, having the highest count of total and unique knowledge for both types. To assess the quality of the distilled knowledge, we recruit four expert annotators to conduct human evaluations on the plausibility of the generated conceptualizations and instantiations. They are asked to annotate the plausibility of 3,000 randomly sampled abstract commonsense triples and 3,000 instantiated triples from the distilled knowledge set. Accepted triples are those deemed plausible by all annotators. We then analyze accepted triple ratios for different levels of critic filtering, as shown in Table 1. Our findings show that LLMs have impressive conceptualization and instantiation abilities, with initial plausibility rates of 82.6% and 77.9% for both types of knowledge, respectively. Critic filtering improves plausibility by up to 14.6% and 16.6%, demonstrating the effectiveness of our measures in maintaining high-quality distilled knowledge. For more annotation details, refer to Appendix E.
@@
The process of abstracting an event into highly diverse conceptualizations plays a crucial role in CANDLE. It is of significant importance because the greater the diversity of conceptualizations, the broader the knowledge coverage becomes upon instantiation. This, in turn, enhances the overall knowledge coverage within the distillation process. To examine the diversity of the top 10,000 popular distilled conceptualizations, we obtain their hypernyms by matching against Probase Wu et al. (2012) and present a visualization in Figure 3. It reveals that our distilled conceptualizations possess a high level of diversity across various categories, forming a comprehensive and intricate knowledge base. Novelty of our distilled conceptualizations and instantiations, compared to other available resources, are in Appendix A.
@@
| CANDLE Distilled (Ours) | RoBERTa-large 340M | 80.69↑2.18 | 80.99↑2.46 | 83.11↑0.84 | 84.50↑1.48 |
| DeBERTa-v3-large 435M | 80.97↑1.42 | 81.14↑1.75 | 83.64↑0.76 | 84.64↑1.12 | |
| LLAMA2 (LoRA Fine-tuned) 7B | 77.48↑1.68 | 78.27↑2.00 | 81.68↑1.79 | 83.40↑1.25 | |
| Mistral-v0.1 (LoRA Fine-tuned) 7B | 77.77↑2.06 | 78.29↑1.53 | 81.95↑2.36 | 82.54↑2.19 | |
| VERA-T5 (Fine-tuned) 5B | 77.54↑1.85 | 78.03↑1.82 | 82.79↑2.66 | 83.61↑2.36 | |